HUBBLE SPACE TELESCOPE DAILY REPORT #2196 PERIOD COVERED: 0000Z (UTC) 08/26/98 - 0000Z (UTC) 08/27/98 Daily Status Report as of 239/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/2 6767 (A Search for Brown Dwarfs and Luminous Young Planets in the Hyades) The Near Infrared Camera (Camera Two) was used to make observations of the brown dwarf star RHY067 which is located in the Hyades Cluster. This is to search for substellar companions orbiting low-mass Hyades M class dwarf stars. Differential imaging was used using methane and water band (on/off) filter pairs. The observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.5 Completed Five Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations take in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 7929 (Cycle 7 WF/PC-2 CTE Monitor) The WF/PC-2 was used to make observations of NGC-5139 to make calibrations observations to monitor the CTE. This Cycle 7 proposal is executed every six months to monitor changes. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed Three Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.8 Completed NIC/1/2/3 7855 (The Evolution of Powerful Radio Galaxies) The Near Infrared Camera (Camera One, Two and Three) was used to make observations of the powerful radio galaxy 3C19. Our proposed NICMOS observations combined with our multi-wavelength observing program will finally provide us with the data we need to build a picture of how powerful radio galaxies evolve. The observations were completed as planned, and no problems were reported. 1.9 Completed STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.10 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.11 Completed WF/PC-2 7476 (Sakurai's Nova-Like Object: Real-Time Monitoring of a Stellar Thermal Pulse) The WF/PC-2 was used to make images of Sakurai's nova-like object in order to do real-time monitoring of a stellar thermal pulse. The eruption of Sakurai's nova-like variable in Sagittarius presents a "once-in-a-lifetime" opportunity for real-time observations of a star undergoing a final helium thermal pulse. The observations were completed as planned, and no anomalies were reported. 1.12 Completed NIC/1/2/3 7886 (NICMOS Snap Shot Survey of Early-Type Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the early-type galaxy Abell2052 as part of a snap-shot broad-band imaging survey. Since almost all early-type galaxies harbor dust, these images which are less sensitive to extinction and will provide superior light profiles for dynamical studies of the cores of these galaxies. The observations were completed as scheduled, and no problems were reported. 1.13 Completed NIC/1 7867 (NICMOS Imaging of the Dusty CfA Seyfert Nuclei) FGS 7968 () The Near Infrared Camera (Camera One) was used to make imaging observations of the dusty CfA class Seyfert galaxy NGC-1144. These data will allow us to determine if the Seyfert 1/2 dichotomy is best explained by the unified model, the evolutionary picture, or a combination of both of these scenarios. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2/STIS/CCD/NIC/1/2/3 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the rich star clusters which are located within the Large Magellanic Cloud in order to study their formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. The observations were completed as scheduled, and no anomalies were reported. 1.15 Completed NIC/1 7833 (Multiplicity Survey of Alpha Persei: Studying the Evolution and Effects of Companions) The Near Infrared Camera (Camera One) was used to make observations of the star Alpha-Persei. Extensive observations have established that binaries are twice as common amongst pre-main sequence stars than older solar neighborhood G-dwarfs. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 24 Successful : 24 Scheduled Reacquisitions : 2 Successful : 2 2.2 FHST Updates: Scheduled : 26 Successful : 25 The roll update at 239/0218Z failed. Two subsequent full maneuver updates were successful, however, and a map taken 10 minutes earlier was acceptable. HSTAR 6638 was written. 2.3 Operations Notes: The SSR EDAC error counter was cleared four times, per ROP SR-1. Another GRGT IP command test was uneventfully completed, with commanding and telemetry receipt at various rates accomplished via both LGA and HGA. This session ran from 238/2147Z until 238/2207Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH